Two spheres of masses 6g nd 8g r moving towards each other in the same line with velocity 10cm/s and 6cm/s respectively.. Find their velocities after collision if the coefficient of restitution is 0.75

Grade:11

2 Answers

Vikas TU
4 years ago
Hiii Dhruv 
Cofficient of restitution = v2 –  v1 / u1 – u2 
Where u1 and u2 are velocity before collision and v1 and v2 are velocity after collision.
so 0.75 = v2 –  v1 / u1 – u2
v2 – v1 = 3............................................(1)
Apply conservation of momentum 
m1u1 + m2u2 = m1v1 + m2v2
so, putting the value in this we get :
6v1 + 8 v2 = 108.................................(2)
Solving equation 1 and 2  we get 
v1 = 6 cm/s and v2 = 9 cm/s 
Khimraj
4 years ago
m1u1 + m2u2 = m1v1 + m2v2
6v1 + 8 v2 = 108.
.........................................................................................